OpenStudy (anonymous):

Help Please! Reduce rational expression, x+3 x to the 2 power-3x-4 ____________________ . ________________________ x to the 2 power-16 x to the 2 power -1 Thank you.

OpenStudy (saifoo.khan):

\[\frac{x+3}{x^2-16} \times \frac{x^2-3x-4}{x^2-1}\]

OpenStudy (saifoo.khan):

\[\frac{x+3}{x^2+3x-4}\]

OpenStudy (anonymous):

ty :)

OpenStudy (saifoo.khan):

You welcome!! :) do u have like a million questions?

OpenStudy (anonymous):

yes!!! so sorry, i'm doing h/w. not very confident in my answers.

OpenStudy (saifoo.khan):

No problem!! lol. it's been like hours you are doing them.

OpenStudy (anonymous):

i got the 3x/ but the bottom i couldn't figure out, do have an explanation. I appreciate your help

OpenStudy (anonymous):

or correction 3+x/

OpenStudy (saifoo.khan):

3+x and 3+x mean the same thing. and you welcome.

OpenStudy (saifoo.khan):

is that college math?

OpenStudy (anonymous):

yes, intermediate algebra

OpenStudy (saifoo.khan):

mathlab?

OpenStudy (anonymous):

is it good...? I'm doing the work i'm just not confident in my answers. I'm not sure what to do, think i need a tutor

OpenStudy (saifoo.khan):

That's a no problem. i have dont these questions like million times that's why i know. i help many people who are stuck with algebra.

OpenStudy (anonymous):

i really appreciate it, sorry if i'm being a pain :(...

OpenStudy (saifoo.khan):

No problem! ^_^

OpenStudy (saifoo.khan):

i can understand that.

OpenStudy (anonymous):

x+3x2−16×x2−3x−4x2−1 (x+3)(x-4)(x+1)/(x+4)(x-4)(x+1)(x-1) (x+3)/(x+4)(x-1)

OpenStudy (saifoo.khan):

in which week are u in? if u want help with your algebra i can help u out as well. :D

OpenStudy (anonymous):

going on my 3rd week. I have a test next monday.

OpenStudy (anonymous):

GT, thank you

OpenStudy (saifoo.khan):

Sure, if u want.. u may email me on sunday.. i can help u with checkpoints and quizes and stuff.

OpenStudy (anonymous):

that would be great. i have review worksheet maybe you could help with me.

OpenStudy (saifoo.khan):

Alright. when is it due?

OpenStudy (anonymous):

monday, right before the test

OpenStudy (saifoo.khan):

i have a final due this sunday. do u have facebook? it will be more convinent. otherwise, let me know on fri, sat. we will work on it as well. how many questions does it has?

OpenStudy (anonymous):

20 questions, my email is cierranicoie@aol.com

OpenStudy (saifoo.khan):

i emailed u.

OpenStudy (anonymous):

got it

OpenStudy (saifoo.khan):

Alright. cool.

OpenStudy (anonymous):

thank you

OpenStudy (saifoo.khan):

No problem. do u have any more problems for today?

OpenStudy (anonymous):

i do actually...